how to calculate degree of hardness of water?
please send as early as possible

Grade:11

5 Answers

Umakant biswal
7 years ago
@ chakri 
we cannot have any special formula for that , but we can find by the following as example 
we need to calculate the hardness if 9.6 mg of Mgso4 present in 1 kg of water , 
so, here 
given , 1000g of water contains Mgso4 = 9.6 mg 
1000000g of water contain mgso4 = 9.6 * 1000 = 9600 mg = 9.6 g 
u need to remember – 100000g = one million parts of water 
1 mole of mgso4 = 1 mole of caco3 
120 g of mgso4 = 100g of caco3 
9.6 g of mgso4 = (100g * 9.6 g ) / 120 g of caco3 = 8g of caco3 
so, here degree of hardness = 8 ppm 
HOPE IT CLEARS 
vishnu budarapu
7 years ago
degree of hardness =(GMwt. of CaCO3 /GMwt. of given salt) X (wt of given salt / wt of taken water) X 1000000
saurabh singh
6 years ago
degree of hardness=strength of hardness producing substance×multiplication factor so,multiplication factor=100/2×chemical equivalent of hardness producing substance.
Om agrawal
3 years ago
We can find hardness of substance woth relative to CaCO3 
Ex- hardness of mgso4 of 90 gram
120 gram MSO4=100 gram CaCO3 
90 gram = 100*90/120
 
75 ppm [ ans]
Yash Chourasiya
3 years ago
Dear Students

we cannot have any special formula for that , but we can find by the following as example
we need to calculate the hardness if 9.6 mg of Mgso4 present in 1 kg of water ,
so, here
given ,
1000g of water contains Mgso4 = 9.6 mg
1000000g of water contain mgso4 = 9.6 * 1000 = 9600 mg = 9.6 g
You need to remember : – 100000g = one million parts of water
1 mole of mgso4 = 1 mole of caco3
120 g of mgso4 = 100g of caco3
9.6 g of mgso4 = (100g * 9.6 g ) / 120 g of caco3 = 8g of caco3
so, here degree of hardness = 8 ppm

I hope this answer will help you.
Thanks & Regards
Yash Chourasiya